# Polynomial-Time Preparation of Low-Temperature Gibbs States for 2D Toric Code

### v1

@@ 4.3 Proof of Lemmas
Appendix A Low temperature Gibbs state preparation for 1D ferromagnetic Ising model
@@
This section studies the low-temperature thermal state preparation for a 1D ferromagnetic Ising chain with periodic boundary condition. Although the Hamiltonian of this model is classical (a diagonal matrix in the computational basis), the jump operators are quantum and involve a significant amount of non-diagonal elements. The analysis of this model also shows the generality of core techniques for the 2D toric code in Section 4.
@@
$H^{\rm Ising}=-J\sum_{j=1}^{N}\sigma^{z}_{j}\sigma^{z}_{j+1}\,,\quad J>0\,,$ (A.1)
@@
$\mathcal{L}_{\beta}=\sum_{j=1}^{N}\left(\mathcal{L}_{\sigma^{x}_{j}}+\mathcal{L}_{\sigma^{y}_{j}}+\mathcal{L}_{\sigma^{z}_{j}}\right)+\mathcal{L}_{\mathsf{X}}\,,$ (A.3)
which is the sum of local Lindbladians with Pauli couplings plus a global one with coupling operator $\mathsf{X}:=\prod_{j=1}^{N}\sigma_{j}^{x}$. Here $\mathcal{L}_{\sigma_{j}^{x/y/z}}$ and $\mathcal{L}_{\mathsf{X}}$ are defined via (2.4). Their explicit formulas can be similarly derived as in (4.16), (4.18), and (4.19). The main result of this section is the following spectral gap lower bound of $\mathcal{L}_{\beta}$ in Eq. A.3.
@@
For the Davies generator (A.3), we have
$\mathrm{Gap}(\mathcal{L}_{\beta})\geq\max\left\{\Theta(e^{-4\beta J}),\Theta(N^{-3})\right\}\,.$
@@
$\mathcal{B}(\mathcal{H})=(I\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{X}\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{Z}\otimes\mathcal{A}^{\rm full}_{b})\,.$ (A.4)
We are now ready to sketch the proof of 17. For simplicity, we use $b_{j}=\pm$ for the bond configuration. Similarly to Lemma 5, the algebra $\mathcal{A}_{b}^{\rm full}$ is generated by the observables $\{\textbf{Z}_{j}\}^{N}_{j=1}$ and $\{\sigma^{x}_{j}\}_{j=2}^{N}$, which commute with $\mathsf{X}$. Then, by a direct computation, we have $\mathcal{L}_{\mathsf{X}}\left(I\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=0$. Moreover, there holds
$\mathcal{L}_{\mathsf{X}}\left(\mathsf{Z}/\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=-2(\mathsf{Z}/\mathsf{Y})\otimes\mathcal{A}^{\rm full}_{\rm b},\quad-\langle(\mathsf{Z}/\mathsf{Y}),\mathcal{L}_{\mathsf{X}}(\mathsf{Z}/\mathsf{Y})\rangle_{\sigma_{\beta}}=2\,.$ (A.5)
@@
$\mathcal{L}_{\mathsf{Z}}\left(\mathsf{X}/\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=-2\left(\mathsf{X}/\mathsf{Y}\right)\otimes\mathcal{A}^{\rm full}_{\rm b}\,,\quad-\langle(\mathsf{X}/\mathsf{Y}),\mathcal{L}_{\mathsf{Z}}(\mathsf{X}/\mathsf{Y})\rangle_{\sigma_{\beta}}=2\,.$ (A.6)
Further, for the action of local Lindbladian $\mathcal{L}_{\sigma^{x}_{j}/\sigma^{y}_{j}/\sigma^{z}_{j}}$ on the decomposition (A.4), we have the following lemma, in analog with Lemma 6.
@@
The Lindbladian with Pauli coupling $\mathcal{L}_{\sigma^{x}_{j}}$, $\mathcal{L}_{\sigma^{y}_{j}}$, and $\mathcal{L}_{\sigma^{z}_{j}}$ ($j\geq 1$) are block diagonal with respect to the decomposition (A.4):
@@
~$\widetilde{\mathcal{L}}=\sum_{j=2}^{N}\mathcal{L}_{\sigma^{x}_{j}}$ (A.7)
which is primitive when restricted on $\mathcal{B}(\mathcal{H}_{+})=\mathcal{A}^{\rm full}_{\rm b}$ [2, Lemma 6]. Thanks to the properties (A.5)–(A.6) of $\mathcal{L}_{\mathsf{X}}$ and $\mathcal{L}_{\mathsf{Z}}$, the Lindbladian ~$\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}$ (as a part of $\mathcal{L}_{\beta}$ in (A.3)) is primitive on $\mathcal{B}(\mathcal{H})$. Then, Lemma 3 (item 2) readily gives
$\mathrm{Gap}\left(\mathcal{L}_{\beta}\right)\geq\mathrm{Gap}\left(\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}\right)\,.$ (A.8)
Thus, it suffices to consider the spectral gap of the latter one. For this, we note from (A.5)–(A.6) that $\mathcal{L}_{\mathsf{X}/\mathsf{Z}}$ is also block diagonal for the decomposition (A.4). Then, by Lemma 19, we only need to estimate the gap of ~$\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}$ on each invariant subspace $I/\mathsf{X}/\mathsf{Y}/\mathsf{Z}\otimes\mathcal{A}_{\rm b}^{\rm full}$:
@@
Combining the above arguments with (A.8), we find

### v2

@@ 4.3 Proof of Lemmas
Proof of Lemma 13.
From (4.39), the matrix representation of the master Hamiltonian of $-\mathcal{L}_{1}$ can be written as:
$K_{L^{2}}^{\beta}=K_{L^{2}}+(K_{L^{2}}^{\beta}-K_{L^{2}})\,,$
with
and $K_{L^{2}}^{\beta}-K_{L^{2}}$ is given by
$\left\lVert K_{L^{2}}^{\beta}-K_{L^{2}}\right\rVert=\Theta(L^{2}e^{-2\beta})=\mathcal{O}(L^{-6})\,,\quad\text{when $\beta\geq 4\ln L$\,.}$
Therefore, to obtain the desired gap estimate (B.9), we only need to consider the gap at zero temperature and prove $\mathrm{Gap}(K_{L^{2}})=\Omega(L^{-4})$.
For this, we consider the following configuration space
[…]
$K_{L^{2}}:\mathcal{S}_{L^{2},k}\to\mathcal{S}_{L^{2},k}\,,$ (4.45)
$K_{L^{2}}=K_{L^{2}-1}\otimes I_{1}+I_{L^{2}-2}\otimes(k_{T}+k_{D})\,.$
In addition, there holds
$a^{*}K_{L^{2}}a$ $\geq a^{*}(K_{L^{2}-1}\otimes I_{1})a$ (4.46)
$=(a_{+})^{*}K_{L^{2}-1}a_{+}+(a_{-})^{*}K_{L^{2}-1}a_{-}$
$\geq\min\left\{\lambda_{L^{2}-1,k},\lambda_{L^{2}-1,k-1}\right\}\,,$
$K_{L^{2}}|_{\mathcal{S}_{L^{2},2}}=(K_{T})|_{\mathcal{S}_{L^{2},2}}+(K_{D})|_{\mathcal{S}_{L^{2},2}}\,.$
Note that the basis of the subspace ${\cal S}_{L^{2},2}$ can be denoted by:
$\left\{|i,j\rangle:1\leq i\leq L^{2}-1,i+1\leq j\leq L^{2}\right\}\,.$
This represents the positions of the two “--” signs and defines a stair graph $G_{L^{2}-1}$ with parameter $L^{2}-1$ (see 17 and Fig. 7 in Appendix A). From the representation (4.42) and (4.43), there holds
$K_{T}|i,j\rangle=\textbf{1}_{i>1}(|i,j\rangle-|i-1,j\rangle)+\textbf{1}_{i<j-1}(|i,j\rangle-|i+1,j\rangle)$
$+\textbf{1}_{j>i+1}(|i,j\rangle-|i,j-1\rangle)+\textbf{1}_{j<L^{2}}(|i,j\rangle-|i,j+1\rangle)\,,$
where the first and third term represent the transition from $\ket{+-}$ to $\ket{-+}$ and the second and fourth term represent the transition from $\ket{-+}$ to $\ket{+-}$. It is straightforward to see that $(K_{T})|_{{\cal S}_{L^{2},2}}$ is exactly the graph Laplacian associated with the stair graph $G_{L^{2}-1}$. Moreover, $(K_{D})|_{{\cal S}_{L^{2},2}}$ can be identified with the diagonal matrix $D_{L^{2}-1}$ in 20:
$K_{D}|i,j\rangle=\textbf{1}_{j=i+1}\cdot 2|i,i+1\rangle\,.$
Hence, by 20 with $n=L^{2}-1$,
$\lambda_{L^{2},2}=\lambda_{\min}((K_{T})|_{\mathcal{S}_{L^{2},2}}+(K_{D})|_{\mathcal{S}_{L^{2},2}})=\lambda_{\min}(H_{L^{2}-1})=\Omega\left(L^{-4}\right)\,.$ (4.47)
Using Eq. 4.46, it follows that $\min_{3\leq k\leq L^{2}}\lambda_{L^{2},k}=\Omega\left(L^{-4}\right)$. Therefore, we have proved
$\mathrm{Gap}\left(K_{L^{2}}\right)=\Omega\left(L^{-4}\right)=\Omega\left(N^{-2}\right).$
Proof of Lemma 15.
$\left\lVert K_{L^{2}}^{\beta}-K_{L^{2}}\right\rVert=\Theta(L^{2}e^{-2\beta})=\mathcal{O}(L^{-6})\,,\quad\text{for $\beta\geq 4\ln L$}\,,$
with
Then, $K_{L^{2}}$ is block diagonal for (4.48) and
We first consider the subspace $\mathcal{S}_{{L^{2}},2}$, whose basis vectors contain only two ``"``"``-"` ` - " signs. Our strategy for lower bounding $\lambda_{\min}\left(K_{L^{2}}|_{\mathcal{S}_{{L^{2}},2}}\right)$ is to reduce this problem to a straight line case as in Fig. 6. For this, we introduce a set of lines covering all vertices in the comb. More specifically, we may regard the comb as a connected graph (more precisely, a tree) with $L^{2}$ vertices of degree at most 3, each corresponding to a star $\textbf{X}_{s}$ that interacts with the comb. Let ${\bf D}_{\rm comb}$ be the set of degree-one vertices (i.e., end-points) in the comb, and $l_{u,v}$ be the shortest path between the vertices $u$ and $v$ in the comb. Define
For an arbitrary unit vector $\alpha\in\mathcal{S}_{{L^{2}},2}$:
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},2}}p_{a}\ket{a}\,,$
by 16, there exists a path $\tilde{l}$ in ${\bf P_{\rm comb}}$ such that for
it holds that
$\|\alpha_{\tilde{l}}\|_{2}^{2}=\Omega(1/L^{2})\,.$ (4.50)
$(K_{L^{2}})_{\tilde{l}}:=\sum_{e=(u,v)\in\tilde{l}}(k_{T}+k_{D})_{u,v}\,,$
$=$ $\leavevmode\nobreak\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega(L^{-2})\cdot\Omega(L^{-2})=\Omega(L^{-4})}\,,$
where the first step follows from $K_{L^{2}}-(K_{L^{2}})_{\tilde{l}}$ is positive semi-definite, the second step follows from $\langle\alpha-\alpha_{\tilde{l}},(K_{L^{2}})_{\tilde{l}}\alpha_{\tilde{l}}\rangle=0$, the third step follows from $(K_{L^{2}})_{\tilde{l}}$ is equivalent to a 1D chain of length $|\tilde{l}|\leq\ell_{\rm comb}=\mathcal{O}(L)$ and $a_{\tilde{l}}\in\mathcal{S}_{|\tilde{l}|,2}$, and the last step follows from (4.50) and (4.47). Thus, we conclude
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K_{L^{2}}|_{\mathcal{S}_{{L^{2}},2}}=\Omega(L^{-4})=\Omega(N^{-2})}\,.$ (4.52)
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},k}}p_{a}\ket{a}\,,$
by 16, there exists a path $\tilde{l}$ in ${\bf P_{\rm comb}}$ such that for
we have
$\sum_{q=2}^{k}\|\alpha_{\tilde{l},q}\|_{2}^{2}=\Omega(1/L^{2})\,.$ (4.53)
$=$ $\leavevmode\nobreak\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega(L^{-2})\cdot\Omega(L^{-2})=\Omega(L^{-4})}\,,$
where the first step follows from $K_{L^{2}}-(K_{L^{2}})_{\tilde{l}}$ is positive semi-definite, the second step follows from $\alpha_{\tilde{l},q_{1}}^{*}(K_{L^{2}})_{\tilde{l}}\alpha_{\tilde{l},q_{2}}=0$ for $q_{1}\neq q_{2}$, the third step follows from $(K_{L^{2}})_{\tilde{l}}$ is equivalent to a 1D chain of length $|\tilde{l}|\leq\ell_{\rm comb}=\mathcal{O}(L)$, and $\alpha_{\tilde{l},q}$ is in the subspace $\mathcal{S}_{|\tilde{l}|,q}$, the fourth step follows from the recursion relation (4.46) that $\lambda_{\ell_{\rm comb},q}\geq\lambda_{\ell_{\rm comb},2}$, an the last step follow from (4.47) and (4.53). Thus, we have
$K_{L^{2}}|_{\mathcal{S}_{{L^{2}},k}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega(L^{-4})=\Omega(N^{-2})}\,.$
Combining them all together, we now can conclude
$\mathrm{Gap}\left(-\mathcal{L}_{2}|_{\mathcal{A}_{\rm e}^{\rm full}([L^{2}])}\right)\geq\min_{2\leq k\leq L^{2}}\lambda_{\min}\left(K_{L^{2}}|_{\mathcal{S}_{{L^{2}},k}}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{O}(N^{-3})=\Omega(N^{-2})}\,.$
This concludes the proof of Lemma 15. ∎
Fact 16.
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},k}}p_{a}\ket{a}\,,$
$\sum_{\begin{subarray}{c}a\in\mathcal{A}_{L^{2},k}:\\ \#\{s\in\tilde{l}:a_{s}=``-"\}>1\end{subarray}}|p_{a}|^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega(1/L^{2})}\,.$
Proof.
$=$ $\leavevmode\nobreak\ 1\,,$
$\sum_{\begin{subarray}{c}a\in\mathcal{A}_{L^{2},k}:\\ \#\{s\in l:a_{s}=``-"\}>1\end{subarray}}|p_{a}|^{2}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1/L^{2}}\,,$
which proves the proposition. ∎
Data availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Conflict of interest. The authors have no conflicts of interest to declare that are relevant to the content of this article.
Appendix A Spectral graph theory problem
In this appendix, we study the spectral property of the stair graph, which is the grid graph folded along the diagonal. Let $n$ be the side length of the stair. We consider the graph Laplacian of the stair graph of size $\binom{n+1}{2}$, perturbed by adding positive weights to the vertex on the diagonal. We shall prove that the minimum eigenvalue of such class of perturbed Laplacians is $\Theta(n^{-2})$. The result is of independent interest and is applied to the proofs of Lemmas 13 and 15, see (4.47) and (4.52) for detail.
We first introduce the definition of the stair graph, graph Laplacian, and their basic properties.
Definition 17 (Stair graph).
A stair graph $G=(V,E)$ with parameter $n\in\mathbb{N}_{+}$ is an undirected graph with $\binom{n+1}{2}=\frac{n(n+1)}{2}$ vertices. It can be viewed as an $n$-by-$n$ grid folded along the diagonal (see Fig. 7 for an example).
More specifically, let the rows be labeled by $1,2,\dots,n$ (from bottom to top), and the columns be labeled by $12,3,\dots,n+$ 1 (from right to left) [^4] These labels are chosen to align with the notation used in the proof of Lemmas 13 and 15.. Then, each vertex gets a label $(i,j)$ based on its row and column indices.
For a vertex $(i,j)$, it is connected to its nearest-neighbors $(i,j\pm 1)$ and $(i\pm 1,j)$, with open boundary condition.
Figure 7: An example of the stair graph ($n=8$) with the vertex labels. The red edges are the horizontal edges $E_{8,h}$. The blue self-loops are the entries in $D_{8}$ of value $2$.
Definition 18 (Graph Laplacian).
For an undirected graph $G=(V,E)$, its graph Laplacian $\mathcal{L}_{G}$ is defined to be:
${\cal L}_{G}:=D_{G}-A_{G}\,,$
where $D_{G}=\operatorname{diag}(\{\deg(v)\}_{v\in V})$ and $A_{G}$ is the adjacency matrix of $G$.
Proposition 19 (Spectral properties of graph Laplacian).
For any undirected graph $G$, its Laplacian ${\cal L}_{G}$ satisfies the following properties:
  • ${\cal L}_{G}$ is a real-symmetric, positive semi-definite matrix with minimum eigenvalue 00.
  • The number of zero eigenvalues is equal to the number of connected components of $G$.
Now, we are ready to present the main theorem in this section:
Theorem 20.
For any $n\in\mathbb{N}_{+}$, let $G_{n}=(V_{n},E_{n})$ be the stair graph with parameter $n$, and ${\cal L}_{n}$ be its graph Laplacian. Let $D_{n}:V_{n}\times V_{n}\rightarrow\mathbb{R}$ be a diagonal matrix such that $D_{n}((i,i+1),(i,i+1))=2$ and other entries equal to 0. Then
$H_{n}:={\cal L}_{n}+D_{n}\succeq\Omega\left(\frac{1}{n^{2}}\right)\,.$
Proof.
For any vector $f\in\mathbb{R}^{V_{n}}$, we have
$f^{\top}{\cal L}_{n}f=\sum_{(u,v)\in E_{n}}(f(u)-f(v))^{2}\,.$
Thus,
$f^{\top}H_{n}f=\sum_{(u,v)\in E_{n}}(f(u)-f(v))^{2}+\sum_{i=1}^{n}2f(i,i+1)^{2}\,.$
In the second term, we abuse the notation and use $f(i,j)$ to denote $f(v)$ with $v=(i,j)$ for simplicity.
For any off-diagonal vertex $v=(i,j)$ with $i+2\leq j\leq n+1$, consider the horizontal path in the stair graph:
$(i,j)\rightarrow(i,j-1)\rightarrow(i,j-2)\rightarrow\cdots\rightarrow(i,i+1)\,.$
We can apply telescoping summation and Cauchy-Schwartz inequality along this path:
$f(i,j)=f(i,i+1)+\sum_{k=i+1}^{j-1}(f(i,k+1)-f(i,k))\,,$
which implies that
$f(i,j)^{2}\leq$ $\leavevmode\nobreak\ 2f(i,i+1)^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\left(\sum_{k=i+1}^{j-1}(f(i,k+1)-f(i,k))\right)^{2}$ ($(a+b)^{2}\leq 2a^{2}+2b^{2}$)
$\leq$ $\leavevmode\nobreak\ 2f(i,i+1)^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}(j-i-1)\cdot\sum_{k=i+1}^{j-1}(f(i,k+1)-f(i,k))^{2}\,.$ (Cauchy-Schwartz)
Next, we sum over all vertices:
$\sum_{v\in V_{n}}f(v)^{2}=$ $\leavevmode\nobreak\ \sum_{i=1}^{n}f(i,i+1)^{2}+\sum_{i=1}^{n-1}\sum_{j=i+2}^{n+1}f(i,j)^{2}$
$\leq$ $\leavevmode\nobreak\ \sum_{i=1}^{n}f(i,i+1)^{2}+\sum_{i=1}^{n-1}\sum_{j=i+2}^{n+1}\left(2f(i,i+1)^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}(j-i-1)\cdot\sum_{k=i+1}^{j-1}(f(i,k+1)-f(i,k))^{2}\right)$
$=$ $\leavevmode\nobreak\ \sum_{i=1}^{n}(2(n-i)+1)f(i,i+1)^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\sum_{i=1}^{n-1}\sum_{k=i+1}^{n}(f(i,k+1)-f(i,k))^{2}\sum_{j=k+1}^{n+1}(j-i-1)$
$=$ $\leavevmode\nobreak\ \sum_{i=1}^{n}(2(n-i)+1)f(i,i+1)^{2}+\sum_{i=1}^{n-1}\sum_{k=i+1}^{n}(f(i,k+1)-f(i,k))^{2}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(n+k-2i)(n-k+1)}$
$\leq$ $\leavevmode\nobreak\ \mathcal{O}(n)\cdot\sum_{i=1}^{n}2f(i,i+1)^{2}+\mathcal{O}(n^{2})\cdot\sum_{(u,v)\in E_{n,h}}(f(u)-f(v))^{2}\,,$
where $E_{n,h}$ are the set of horizontal edges in $G_{n}$, and the last step follows from $2(n-i)+1\leq 2n$ and $(n+k-2i)(n-k+1)\leq n(n-1)$.
Since $E_{n,h}\subset E_{n}$, we have
$\|f\|_{2}^{2}=\sum_{v\in V_{n}}f(v)^{2}\leq$ $\leavevmode\nobreak\ \mathcal{O}(n^{2})\cdot\left(\sum_{i=1}^{n}2f(i,i+1)^{2}+\sum_{(u,v)\in E_{n}}(f(u)-f(v))^{2}\right)$
$=$ $\leavevmode\nobreak\ \mathcal{O}(n^{2})\cdot f^{\top}H_{n}f\,.$
By the variational characterization of eigenvalues,
$\lambda_{\min}(H_{n})=\min_{f\in\mathbb{R}^{V_{n}}}\frac{f^{\top}H_{n}f}{\|f\|_{2}^{2}}=\Omega\left(\frac{1}{n^{2}}\right)\,.$
The following lemma shows that the $1/{n^{2}}$ lower bound is tight:
Lemma 21.
For $n\in\mathbb{N}_{+}$, let $H_{n}$ be defined as in 20. Then,
$\lambda_{\min}(H_{n})\leq\mathcal{O}\left(\frac{1}{n^{2}}\right)\,.$
Proof.
We can construct a test vector $g\in\mathbb{R}^{V_{n}}$ such that:
$g(i,j):=j-i-1\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall 1\leq i\leq n,\,i+1\leq j\leq n+1\,.$
Then, $g(i,i+1)=0$ for all $i\in[n]$. And for each edge $(u,v)\in E_{n}$, it is easy to check that
$(g(u)-g(v))^{2}=1\,.$
Thus,
$g^{\top}H_{n}g=0+\sum_{(u,v)\in E_{n}}(g(u)-g(v))^{2}=|E_{n}|=n(n-1)\,.$
On the other hand,
$\|g\|_{2}^{2}=\sum_{i=1}^{n}\sum_{j=i+1}^{n+1}(j-i-1)^{2}=\sum_{i=1}^{n}\sum_{k=1}^{n-i}k^{2}=\frac{1}{12}n^{2}(n^{2}-1)\,.$
Thus, by Rayleigh quotient,
$\lambda_{\min}(H_{n})\leq\frac{g^{\top}H_{n}g}{\|g\|_{2}^{2}}=\mathcal{O}\left(\frac{1}{n^{2}}\right)\,.$
Therefore, we conclude that $\lambda_{\min}(H_{n})=\Theta(n^{-2})$.
Remark 22.
20 and Lemma 21 can be easily generalized to the setting where
$H_{n}(a):=\mathcal{L}_{n}+aD_{n}\,,$
where $a=\Omega(1/n)$. In this case, the proof of 20 still implies that the minimum eigenvalue $\lambda_{\min}(H_{n}(a))=\Omega(1/n^{2})$, and the upper bound construction in Lemma 21 does not depend on $a$. Therefore, we have $\lambda_{\min}(H_{n}(a))=\Theta(1/n^{2})$.
@@ Appendix A Spectral graph theory problem
Appendix B Low temperature Gibbs state preparation for 1D ferromagnetic Ising model
@@
This section studies the low-temperature thermal state preparation for a 1D ferromagnetic Ising chain with the periodic boundary condition. Although the Hamiltonian of this model is classical (a diagonal matrix in the computational basis), the jump operators are quantum and involve a significant amount of non-diagonal elements. The analysis of this model also shows the generality of core techniques for the 2D toric code in Section 4.
@@
$H^{\rm Ising}=-J\sum_{j=1}^{N}\sigma^{z}_{j}\sigma^{z}_{j+1}\,,\quad J>0\,,$ (B.1)
@@
$\mathcal{L}_{\beta}=\sum_{j=1}^{N}\left(\mathcal{L}_{\sigma^{x}_{j}}+\mathcal{L}_{\sigma^{y}_{j}}+\mathcal{L}_{\sigma^{z}_{j}}\right)+\mathcal{L}_{\mathsf{X}}\,,$ (B.3)
which is the sum of local Lindbladians with Pauli couplings plus a global one with coupling operator $\mathsf{X}:=\prod_{j=1}^{N}\sigma_{j}^{x}$. Here $\mathcal{L}_{\sigma_{j}^{x/y/z}}$ and $\mathcal{L}_{\mathsf{X}}$ are defined via (3.4). Their explicit formulas can be similarly derived as in (4.16), (4.18), and (4.19). The main result of this section is the following spectral gap lower bound of $\mathcal{L}_{\beta}$ in Eq. B.3.
@@
For the Davies generator (B.3), we have
$\mathrm{Gap}(\mathcal{L}_{\beta})\geq\max\left\{\Theta(e^{-4\beta J}),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Theta(N^{-2})}\right\}\,.$
@@
$\mathcal{B}(\mathcal{H})=(I\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{X}\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{b})\oplus(\mathsf{Z}\otimes\mathcal{A}^{\rm full}_{b})\,.$ (B.4)
We are now ready to sketch the proof of 23. For simplicity, we use $b_{j}=\pm$ for the bond configuration. Similarly to Lemma 5, the algebra $\mathcal{A}_{b}^{\rm full}$ is generated by the observables $\{\textbf{Z}_{j}\}^{N}_{j=1}$ and $\{\sigma^{x}_{j}\}_{j=2}^{N}$, which commute with $\mathsf{X}$. Then, by a direct computation, we have $\mathcal{L}_{\mathsf{X}}\left(I\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=0$. Moreover, there holds
$\mathcal{L}_{\mathsf{X}}\left(\mathsf{Z}/\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=-2(\mathsf{Z}/\mathsf{Y})\otimes\mathcal{A}^{\rm full}_{\rm b},\quad-\langle(\mathsf{Z}/\mathsf{Y}),\mathcal{L}_{\mathsf{X}}(\mathsf{Z}/\mathsf{Y})\rangle_{\sigma_{\beta}}=2\,.$ (B.5)
@@
$\mathcal{L}_{\mathsf{Z}}\left(\mathsf{X}/\mathsf{Y}\otimes\mathcal{A}^{\rm full}_{\rm b}\right)=-2\left(\mathsf{X}/\mathsf{Y}\right)\otimes\mathcal{A}^{\rm full}_{\rm b}\,,\quad-\langle(\mathsf{X}/\mathsf{Y}),\mathcal{L}_{\mathsf{Z}}(\mathsf{X}/\mathsf{Y})\rangle_{\sigma_{\beta}}=2\,.$ (B.6)
Further, for the action of local Lindbladian $\mathcal{L}_{\sigma^{x}_{j}/\sigma^{y}_{j}/\sigma^{z}_{j}}$ on the decomposition (B.4), we have the following lemma, in analog with Lemma 6.
@@
The Lindbladian with Pauli coupling $\mathcal{L}_{\sigma^{x}_{j}}$, $\mathcal{L}_{\sigma^{y}_{j}}$, and $\mathcal{L}_{\sigma^{z}_{j}}$ ($j\geq 1$) are block diagonal with respect to the decomposition (B.4):
@@
~$\widetilde{\mathcal{L}}=\sum_{j=2}^{N}\mathcal{L}_{\sigma^{x}_{j}}$ (B.7)
which is primitive when restricted on $\mathcal{B}(\mathcal{H}_{+})=\mathcal{A}^{\rm full}_{\rm b}$ [2, Lemma 6]. Thanks to the properties (B.5)–(B.6) of $\mathcal{L}_{\mathsf{X}}$ and $\mathcal{L}_{\mathsf{Z}}$, the Lindbladian ~$\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}$ (as a part of $\mathcal{L}_{\beta}$ in (B.3)) is primitive on $\mathcal{B}(\mathcal{H})$. Then, Lemma 3 (item 2) readily gives
$\mathrm{Gap}\left(\mathcal{L}_{\beta}\right)\geq\mathrm{Gap}\left(\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}\right)\,.$ (B.8)
Thus, it suffices to consider the spectral gap of the latter one. For this, we note from (B.5)–(B.6) that $\mathcal{L}_{\mathsf{X}/\mathsf{Z}}$ is also block diagonal for the decomposition (B.4). Then, by Lemma 25, we only need to estimate the gap of ~$\widetilde{\mathcal{L}}+\mathcal{L}_{\mathsf{X}}+\mathcal{L}_{\mathsf{Z}}$ on each invariant subspace $I/\mathsf{X}/\mathsf{Y}/\mathsf{Z}\otimes\mathcal{A}_{\rm b}^{\rm full}$:
@@
Combining the above arguments with (B.8), we find